# A Deep Dive Into How Open-Source Project Maintainers Review and Resolve Bug Bounty Reports

### v1

@@ 1 Introduction
Prior work has investigated the benefits [47, 15, 20, 59, 38, 17, 69, 45] and challenges [64, 57, 56, 71] of maintaining a bug bounty program, but there does not exist any work that studies bug bounty programs and platforms being applied to the OSS ecosystem. While several studies focus on the perspectives of bug bounty hunters [13, 14, 36, 46, 25, 68, 42] and the security practices of OSS contributors [34, 66], there exist knowledge gaps in the connection between bug bounty programs and OSS projects concerning the perspectives of OSS maintainers. This particular perspective is important since OSS project maintainers typically do not have a security background or adequate funding [62], and vulnerability disclosure is a key component of further securing the OSS ecosystem.
@@
We present results in Section 5. In summary, OSS project maintainers find private disclosure and project visibility to be the most beneficial aspects of bug bounty report review, while hunters’ focus on money or CVEs, as opposed to improving software security, and pressure to review bug bounty reports to be the most challenging. We also find that the most helpful features are incentivized reporting, e.g., bug hunters receiving points for reporting valid vulnerabilities, and security metric calculation assistance. Further, we find collaboration features, e.g., the ability to request assistance from security experts, and more hunter reputation features are the most wanted capabilities on bug bounty platforms.
@@ 2 Related Work
We first sought to understand the benefits of bug bounty programs, their challenges, and how they are evaluated. In particular, we want to know what aspects of successful or unsuccessful bug bounty programs stand out the most.
Several studies highlight the benefits of maintaining a bug bounty program, which includes their role in creating an improved security posture within an organization [47, 15, 20, 71], responsible vulnerability disclosure [59], and increased number of reported vulnerabilities [38, 17, 69]. Walshe & Simpson demonstrate that, on average, maintaining a bug bounty program has a comparatively lower yearly cost than hiring two additional software engineers [65]. How to successfully incentivize hackers while keeping total costs low and avoiding a diverted focus from security to financial incentives has also been studied through a game model [35] and analyzing participation and payout rates [26]. Having a bug bounty program also comes with challenges, primarily dealing with invalid, low-quality reports from bug bounty hunters [64, 57, 71]. However, work exists centered around mitigating such issues [56, 49].
@@
We then looked at work that focused on the perspectives of bug bounty hunters, who have also been referred to as white-hat hackers and security researchers throughout the literature. Akgul et al. conducted various surveys and interviews to understand the bug bounty hunters’ motivations and challenges when participating in the bug bounty ecosystem [13]. They found that rewards and learning opportunities drive them, while communication problems demotivate them the most. Several other studies have investigated how bug bounty hunters operate, particularly what incentivizes them to participate in bug bounty programs [14, 36, 46]. Two studies have investigated how to best collaborate with bug bounty hunters to achieve effective program performance through the perspective of control theory and relations governance mechanisms [68, 42]. Ellis & Stevens provide significant socio-technical insights into bug bounty hunters, including discussions on the risks and insecurities they face as essentially gig workers [25].
Our understanding of OSS project contributors is rooted in work that has investigated their existing security practices and culture. Hendrick and Ramaswami surveyed 441 OSS contributors, 90 of whom identified as maintainers, to gather insights on how they follow and incorporate secure software development best practices in their projects [34]. Unlike contributors who only make occasional contributions, maintainers are deeply familiar with their OSS projects and actively shape their future releases. Some key takeaways from their findings of the maintainers’ perspectives are that 69% of OSS contributors desire clearly defined secure software development best practices, 27% of maintainers define their project’s OSS security policy, and 30% of maintainers implement such policy. Wermke et al. conducted 27 semi-structured interviews with OSS contributors to determine how researchers can better support them in their security and trust practices [66]. One important finding was the need for additional consideration for the diversity of OSS project size, their amount of resources, utilization of security and trust processes, and underlying motivations. While our work indirectly addresses OSS contributors in the form of bug hunters, we specifically focus on maintainers.
@@ 3 Methodology
We adopt the methodology of a recent paper that investigated bug bounty hunter perspectives [13], i.e., those who submit bug bounty reports, to explore the perspectives of OSS project maintainers, i.e., those who review bug bounty reports for OSS projects.
We designed and conducted three studies to investigate our research questions. An initial listing survey study to determine characteristics (RQ1), a Likert-scale rating survey study (RQ2), and an interview study. The first two studies allow us to understand the benefits, challenges, helpful features, and needs of OSS project maintainers. We conduct a follow-up interview study to contextualize results.
All data was found in public spaces, e.g., GitHub profiles. As a result, we skipped recruiting OSS project maintainers if their contact information was not publicly available. Our institution’s ethics review board approved all three studies. Participants signed consent forms detailing study plans and participant rights before data collection. Further, our study is GDPR-compliant.
@@ 3.1 Listing survey study
To identify characteristics, e.g., benefits, when performing bug bounty report review, we conducted an online survey on OSS project maintainers who have validated vulnerabilities using the huntr bug bounty platform ($n_{1}=51$).
@@ 3.1.1 Participant recruitment and piloting
We reached out to OSS project maintainers who have validated vulnerabilities via bug bounty report disclosure using the huntr bug bounty platform from July 2021 to December 2023, covering 558 GitHub projects, and had publicly available contact information in their GitHub security policy or associated with the owning GitHub account.
@@ 3.2 Likert scale survey study
The listing study identified a thorough set of characteristics considered by OSS project maintainers, but not their relative importance. Thus, we conducted a second study asking participants to rate how important each characteristic is to them using a five-point Likert scale ($n_{2}=90$).
@@ 3.2.1 Participant recruitment and piloting
We reached out to OSS project maintainers who have validated vulnerabilities using the huntr bug bounty platform from July 2021 to January 2024, covering 612 GitHub projects, and had publicly available contact information in their GitHub security policy or the owning GitHub account. We also expanded our recruitment to include OSS projects that have used third-party delegated bounty programs, such as Bountysource [10], which was a crowdsourcing website for open-source bounties, and OSS projects that have used bug bounty programs generally, such as HackerOne [3], which also offers a version of their popular HackerOne Bounty program for free to eligible open-source projects; thus, resulting in an additional 509 OSS projects covered. We allowed participants to submit responses during and after conducting interviews in hopes of receiving more responses.
@@ 3.2.2 Survey details
Survey participants were asked to rate each of the 40 characteristics on a five-point Likert scale, which are shown as rows in Table II. We again ended with standard demographic questions to understand our sample population.
@@ 3.3 Interview study
We asked consenting and interested listing study participants to participate in a remote semi-structured interview to learn more about why the identified factors they listed were especially important to them ($n_{3}=17$).
@@ 3.3.1 Participant recruitment and piloting
Interviewees were a subset of the initial Listing survey study described in Section 3.1. We invited all eligible, interested listing study participants to participate in remote semi-structured interviews. 24 participants originally agreed to be interviewed. After reaching out using publicly listed contact information from GitHub projects, 19 responded; afterward, 17 participants showed. We conducted pilots with three participants to test our semi-structured interview protocol. Based on the quality of the interviewees’ responses, we revised our interview questions to further narrow our research focus.
@@ 3.4 Limitations
Our gift card incentive might have been most attractive to interview participants who make less money in OSS maintenance, perhaps due to lower skill or experience; however, OSS project maintainers typically are not funded nor have members that are primarily security-oriented, so our participants are likely to participate for the interest of further securing the OSS ecosystem rather than monetary motivation. We addressed this by recruiting all valid listing and interview studies’ participants directly through email, and Likert-scale study participants from GitHub projects of publicly disclosed bug bounty reports stemming from various bug bounty programs.
@@ 4 Participants
Table I summarizes participants’ self-reported demographics and experiences. We had 51 participants in the listing study, 90 in the Likert-scale rating study, and 17 interviewees. Participants were mainly from Europe and North America and were overwhelmingly male.
@@ 5.1 Vulnerability disclosure
The nature of OSS reflects open vulnerability disclosure after a patch is applied. Although all interviewees agree with this, some have had negative experiences with bug hunters, resulting in early disclosure, and some have neutral or negative feelings towards the idea of disclosure in general. OSS project maintainers also have mixed feelings about the CVE process and bug bounties in general.
@@ 5.1.1 Private disclosure
Interviewees felt that private disclosure allows the necessary time to create a patch before notifying users to upgrade ($I=3$). One interviewee mentioned how they do not mind if a bug hunter opens a GitHub issue [29] to report a vulnerability. In contrast, all other interviewees ($I=16$) encourage bug hunters to reach out via email or other processes specified in security policies that allow private disclosure. In another instance, an interviewee mentioned how their project handles “very vital data” and “being compromised would potentially put my job in a tricky position”, further supporting how even OSS project maintainers emphasize private disclosure before publishing a vulnerability.
@@ 5.1.2 Experiences
Most interviewees do not mention negative experiences with the vulnerability disclosure process ($I=14$). Some believe that the bug bounty process alone should sway hunters from disclosing vulnerabilities early ($I=4$); however, few interviewees experienced early vulnerability disclosure ($I=3$), i.e., the vulnerability was disclosed before approval by the OSS project maintainer. One interviewee describes a dilemma they faced, “we recently had a case where the component they released announced that vulnerability with the steps to follow same day as the patch and it was a day before Christmas holidays,” two others had similar experiences happen on multiple occasions ($I=2$). These disclosures associate negativity with bug bounty programs and may sway maintainers from considering their usage.
Some interviewees describe bug hunters as being “pushy” ($I=3$), and others feel “threatened” ($I=2$) in some cases. In particular, one interviewee described, “If I don’t reply right now, could the reporter be tempted to disclose or exploit the vulnerability they reported?” Additionally, the sense of urgency and pressure raises concerns about the impact on collaboration between bug hunters and project maintainers, who described feeling overwhelmed and having had negative experiences ($I=5$). Furthermore, interviewees argue how they need revised bug hunter reputation systems and guidelines to foster a more cooperative and safe environment during the report review ($I=2$).
@@ 5.1.3 Perspectives
Overall, most interviewees view vulnerability disclosure positively ($I=14$). Some describe disclosing a vulnerability after patching as “one of the strengths of open source” ($I=4$), as it allows dependencies and affected projects to be informed of updates as soon as possible. Further, several interviewees felt that if vulnerabilities were mishandled or it was not clear that a necessary patch was deployed, there would be a negative impact and be reflected upon the project ($I=7$). As one interviewee described, “it is important to users to show that we are listening to bug reports.”
On the other hand, some interviewees argued that they had no choice and might as well keep it transparent to show people the vulnerability was fixed, but believed it looked bad upon the respective project ($I=3$). These same interviewees also reference prior experience using vulnerable dependencies but are unaffected by the specific vulnerability stated. One interviewee stated, ”clearly [this] is stupid because this doesn’t affect the software”, but felt obligated to apply a patch, disclose the vulnerability, and notify users to upgrade.
Interviewees reference CVE creation support as an essential feature ($I=7$) as it streamlined and assisted the creation process during bug bounty report review. Some perceive the CVE process as complicated ($I=6$), where one even described it as a “broken system.” Another interviewee affirms that they “don’t give a shit about the CVE, I care about the problem being solved” and feels that CVEs are a “very weird incentive” for bug hunters.
@@ 5.2.1 Security posture
Improved security posture was the most popular item in our listing study and the fourth-ranked benefit from the Likert-scale study ($L=33$, $\pi=0.124$). Further, project visibility was the third-most listed benefit and ranked second in significance ($L=13$, $\pi=0.142$) respective to other characteristics discovered, reflecting an appreciation towards exposure and openness to criticism.
Interviewees expressed that having “many eyes” on the project and involving the community in searching for vulnerabilities ($I=9$) allows vulnerability detection that is otherwise impossible for OSS project maintainers alone. In particular, one interviewee stated, “When we implement something and maintain it for a number of years, we try to think of all possible edge cases and put ourselves in the shoes of an attacker trying to secure stuff; but it’s inevitable that we miss some things.” Others also mentioned how project exposure allowed older security bugs, some as old as ten years, to be discovered and patched ($I=3$). Interviewees also recognize the importance of fixing vulnerabilities causing a “ripple effect” in the software supply chain ($I=3$), e.g., alerting affected projects to upgrade.
On the other hand, others had received and reviewed reports that were deemed low-quality because of their minor contributions ($I=6$). For instance, one interviewee described how after applying a fix to a low-quality report they received, the “real world quality of the software has essentially not really been improved”, leaving them feeling like their time was wasted and an established negative association with receiving bug bounty report notifications. Results regarding OSS project maintainers’ perspectives on low-quality and spam reports are presented in Section 5.3.1.

### v2

@@ 1 Introduction
Prior work has investigated the benefits [52, 16, 23, 64, 42, 20, 76, 50] and challenges [71, 62, 61, 78] of maintaining a bug bounty program, but there does not exist any work that studies bug bounty programs and platforms being applied to OSS. While several studies focus on the perspectives of bug bounty hunters [13, 14, 39, 51, 28, 75, 46] and the security practices of OSS contributors [37, 73, 17, 18], there exist knowledge gaps in the connection between bug bounty programs and OSS projects concerning the perspectives of OSS maintainers. This particular perspective is important since OSS project maintainers typically do not have a security background or adequate funding [67], and vulnerability disclosure is a key component of further securing the OSS ecosystem.
Privacy and confidentiality of communications, before a patch is disclosed, were highly valued by maintainers in our study, which Alomar et al. [15] do not discuss, as well as the pressure to review reports by maintainers. Further, while Akgul et al. [13] do not emphasize the importance of having proof-of-concept exploits or test cases for the vulnerability, our study highlights that point strongly as the third most challenging aspect in bug bounty report review and in terms of regression testing for CI/CD. However, Votipka et al. [70] point out that bug bounty hunters and testers both say PoCs are needed to communicate the importance and severity of a vulnerability; in particular, disagreements about severity arise from bug bounty hunters when reports might not be reviewed, further annoying the hunters in the process. Understanding how such issues arise is understudied in the context of the OSS ecosystem, which our work investigates.
@@
We present results in Section 5. In summary, OSS project maintainers find private disclosure and project visibility to be the most beneficial aspects of bug bounty report review, while hunters’ focus on money or CVEs, as opposed to improving software security, and pressure to review bug bounty reports to be the most challenging. We also find that the most helpful features are incentivized reporting, e.g., bug hunters receiving points for reporting valid vulnerabilities, and security metric calculation assistance. Further, we find collaboration features, e.g., the ability to request assistance from security experts, and more hunter reputation features to be the most wanted capabilities on bug bounty platforms.
@@ 2 Related Work
Bug Bounty Program Benefits and Challenges. Several studies highlight the benefits of maintaining a bug bounty program, which includes their role in creating an improved security posture within an organization [52, 16, 23, 78], responsible vulnerability disclosure [64], and increased number of reported vulnerabilities [42, 20, 76]. Walshe & Simpson demonstrate that, on average, maintaining a bug bounty program has a comparatively lower yearly cost than hiring two additional software engineers [72]. How to successfully incentivize hackers while keeping total costs low and avoiding a diverted focus from security to financial incentives has also been studied through a game model [38] and analyzing participation and payout rates [29]. Having a bug bounty program also comes with challenges, primarily dealing with invalid, low-quality reports from bug bounty hunters [71, 62, 78]. However, work exists centered around mitigating such issues [61, 54].
@@
Challenges and Dynamics of Bug Bounty Hunters. Akgul et al. conducted various surveys and interviews to understand the bug bounty hunters’ motivations and challenges when participating in the bug bounty ecosystem [13]. They found that rewards and learning opportunities drive them, while communication problems demotivate them the most. Several other studies have investigated how bug bounty hunters operate, particularly what incentivizes them to participate in bug bounty programs [14, 39, 51]. Two studies have investigated how to best collaborate with bug bounty hunters to achieve effective program performance through the perspective of control theory and relations governance mechanisms [75, 46]. Ellis & Stevens provide significant socio-technical insights into bug bounty hunters, including discussions on the risks and insecurities they face as essentially gig workers [28].
Security Practices of OSS Contributors. Our understanding of OSS project contributors is rooted in work that has investigated their existing security practices and culture. Hendrick and Ramaswami surveyed 441 OSS contributors, 90 of whom identified as maintainers, to gather insights on how they follow and incorporate secure software development best practices in their projects [37]. Unlike contributors who only make occasional contributions, maintainers are deeply familiar with their OSS projects and actively shape their future releases. Some key takeaways from their findings of the maintainers’ perspectives are that 69% of OSS contributors desire clearly defined secure software development best practices, 27% of maintainers define their project’s OSS security policy, and 30% of maintainers implement such policy. Wermke et al. conducted 27 semi-structured interviews with OSS contributors to determine how researchers can better support them in their security and trust practices [73]. One important finding was the need for additional consideration for the diversity of OSS project size, their amount of resources, utilization of security and trust processes, and underlying motivations. While our work indirectly addresses OSS contributors in the form of bug hunters, we specifically focus on maintainers.
@@ 3 Methodology
We adopt the methodology of a recent paper that investigated bug bounty hunter perspectives [13], i.e., those who submit bug bounty reports, to explore the perspectives of OSS project maintainers, i.e., those who review bug bounty reports for OSS projects. We designed and conducted three studies to investigate our research questions. An initial listing survey study to determine characteristics (RQ1), a Likert-scale rating survey study (RQ2), and an interview study. The first two studies allow us to understand the benefits, challenges, helpful features, and needs of OSS project maintainers. We conduct a follow-up interview study to contextualize results.
3.1 Ethical considerations
Our institution’s ethics review board approved all three studies. Participants signed consent forms detailing study plans and rights before data collection. Further, our study is GDPR-compliant. We worked with our institution’s IRB to employ any changes that would help reduce complaints or negative feedback from potential subjects. For our recruitment, we identified potential subjects from the huntr platform and manually visited their respective project pages. We refrained from using commits as a source of information for author emails, and instead, skipped over potential subjects if their contact information was not easily accessible, e.g., from their GitHub profile. This approach follows the recommendation Utz et al. [69]: “for future recruitment of study participants we recommend, as also suggested by the DPA, to only use contact information that has visibly been made public by the individuals themselves with the intention of allowing the general public to contact them.”
@@ 3.2 Listing survey study
To identify characteristics, e.g., benefits, when performing bug bounty report review, we conducted an online survey between January 2024 and March 2024 on OSS project maintainers who have validated vulnerabilities using the huntr bug bounty platform ($n_{1}=51$).
@@ 3.2.1 Participant recruitment and piloting
We reached out to OSS project maintainers who have validated vulnerabilities via bug bounty report disclosure using the huntr bug bounty platform from July 2021 to December 2023, covering 558 GitHub projects, and had publicly available contact information in their GitHub security policy or associated with the owning GitHub account. In particular, huntr originally supported any GitHub project. Protect AI acquired huntr in mid-2023 and offboarded projects from the platform that were not AI/ML-related; thus, our targeted subjects cover general OSS projects.
@@ 3.3 Likert scale survey study
The listing study identified a thorough set of characteristics considered by OSS project maintainers, but not their relative importance. Thus, we conducted a second study between February 2024 and May 2024 asking participants to rate how important each characteristic is to them using a five-point Likert scale ($n_{2}=90$).
@@ 3.3.1 Participant recruitment and piloting
We reached out to OSS project maintainers who have validated vulnerabilities using the huntr bug bounty platform from July 2021 to January 2024, covering 612 GitHub projects, and had publicly available contact information in their GitHub security policy or the owning GitHub account. We also expanded our recruitment to include OSS projects that have used third-party delegated bounty programs, such as Bountysource [10], which was a crowdsourcing website for open-source bounties, and OSS projects that have used bug bounty programs generally, such as HackerOne [3], which also offers a version of their popular HackerOne Bounty program for free to eligible open-source projects; thus, resulting in an additional 509 OSS projects covered. We allowed participants to submit responses during and after conducting interviews in hopes of receiving more responses. Further, roughly 20% of Likert survey recruited participants are from the Listing survey, so a majority of the Likert survey participants are likely new participants, thereby minimizing the effect of bias for the Likert survey study.
@@ 3.3.2 Survey details
Survey participants were asked to rate each of the 40 characteristics on a five-point Likert scale, which are shown as rows in Table II. Again, we ended with standard demographic questions to understand our sample population.
@@ 3.4 Interview study
We asked consenting and interested listing study participants to participate in a remote semi-structured interview between February 2024 and May 2024 to learn more about why the identified factors they listed were especially important to them ($n_{3}=17$).
@@ 3.4.1 Participant recruitment and piloting
Interviewees were a subset of the initial Listing survey study described in Section 3.2. We invited all eligible—i.e., we required that (1) the participant is an OSS maintainer and (2) the participant has reviewed more than 2 (median) bug bounty reports because we wanted to capture various reviewing experience levels—and interested listing study participants to participate in remote semi-structured interviews. 24 participants originally agreed to be interviewed. After reaching out using publicly listed contact information from GitHub projects, 19 responded; afterward, 17 participants showed up. We conducted pilots with three participants to test our semi-structured interview protocol. Based on the quality of the interviewees’ responses, we revised our interview questions to further narrow our research focus. In particular, we made minor updates to four interview questions so that they were clearer to reduce the amount of time spent on follow-up clarifications and removed two questions entirely as they were repetitive. Furthermore, we iteratively updated the interview guide based on conducted interviews and participant feedback. Changes were limited to adding a few follow-up questions and minor structural modifications, reaching saturation without any changes past the 11th interview.
@@ 3.5 Limitations
OSS project maintainers typically are not funded nor have members that are primarily security-oriented, so our participants are likely to participate in the interest of further securing the OSS ecosystem rather than monetary motivation, i.e., survey participants are not compensated. We addressed this by recruiting all valid listing and interview studies’ participants directly through email, and Likert-scale study participants from GitHub projects of publicly disclosed reports from various bug bounty programs.
@@ 4 Participants
Table I summarizes participants’ self-reported demographics and experiences. We had 51 participants in the listing study, 90 in the Likert-scale rating study, and 17 interviewees. Participants were mainly from Europe and North America and were overwhelmingly male, which is consistent with OSS developer qualitative studies, where the Americas and Europe comprise most participants [40, 47].
@@ 5.1 Vulnerability disclosure
The nature of OSS reflects open vulnerability disclosure after a patch is applied. Although all interviewees agree with this, some have had negative experiences with bug hunters, resulting in early disclosure, and some have neutral or negative feelings towards the idea of disclosure in general. OSS project maintainers also have mixed feelings about the CVE process and bug bounties in general. For instance, in Alomar et al. [15], a participant mentioned that developers might feel embarrassed or blamed if someone reports a vulnerability in their code, and argued how such attitudes can discourage developers from addressing reports properly.
@@ 5.1.1 Private disclosure
Interviewees felt that private disclosure allows the necessary time to create a patch before notifying users to upgrade ($I=3$). One interviewee mentioned how they do not mind if a bug hunter opens a GitHub issue [32] to report a vulnerability. In contrast, all other interviewees ($I=16$) encourage bug hunters to reach out via email or other processes specified in security policies that allow private disclosure. In another instance, an interviewee mentioned how their project handles “very vital data” and “being compromised would potentially put my job in a tricky position” (P7), further supporting how even OSS project maintainers emphasize private disclosure before publishing a vulnerability.
@@ 5.1.2 Experiences
Most interviewees do not mention negative experiences with the vulnerability disclosure process ($I=14$). Some believe that the bug bounty process alone should sway hunters from disclosing vulnerabilities early ($I=4$); however, few interviewees experienced early vulnerability disclosure ($I=3$), i.e., the vulnerability was disclosed before approval by the OSS project maintainer. One interviewee describes a dilemma they faced, “we recently had a case where the component they released announced that vulnerability with the steps to follow same day as the patch and it was a day before Christmas holidays” (P9), two others had similar experiences happen on multiple occasions ($I=2$). These disclosures associate negativity with bug bounty programs and may sway maintainers from considering their usage.
Some interviewees describe bug hunters as being “pushy” ($I=3$), and others feel “threatened” ($I=2$) in some cases. In particular, one interviewee described, “If I don’t reply right now, could the reporter be tempted to disclose or exploit the vulnerability they reported?” (P2). Additionally, the sense of urgency and pressure raises concerns about the impact on collaboration between bug hunters and project maintainers, who described feeling overwhelmed and having had negative experiences ($I=5$). Furthermore, interviewees argue how they need revised bug hunter reputation systems and guidelines to foster a more cooperative and safe environment during the report review ($I=2$).
@@ 5.1.3 Perspectives
Most interviewees view vulnerability disclosure positively ($I=14$), which is consistent with Alomar et al. [15] participants’ perspectives. Some describe disclosing a vulnerability after patching as “one of the strengths of open source” ($I=4$), as it allows dependencies and affected projects to be informed of updates as soon as possible. Further, several interviewees felt that if vulnerabilities were mishandled or it was not clear that a necessary patch was deployed, there would be a negative impact and be reflected upon the project ($I=7$). As one interviewee described, “it is important to users to show that we are listening” (P7).
On the other hand, some interviewees argued that they had no choice and might as well keep it transparent to show people the vulnerability was fixed, but they believed it looked bad for the respective project ($I=3$). These same interviewees also reference prior experience using vulnerable dependencies but are unaffected by the specific vulnerability stated. One interviewee stated, ”clearly [this] is stupid because this doesn’t affect the software” (P2), but felt obligated to apply a patch, disclose the vulnerability, and notify users to upgrade.
Interviewees reference CVE creation support as an essential feature ($I=7$) as it streamlined and assisted the creation process during bug bounty report review. Some perceive the CVE process as complicated ($I=6$), where one even described it as a “broken system” (P8). Another interviewee affirms that they “don’t give a shit about the CVE, I care about the problem being solved” and feels that CVEs are a “very weird incentive” (P3).
@@ 5.2.1 Security posture
Improved security posture was the most popular item in our listing study and the fourth-ranked benefit from the Likert-scale study ($L=33$, $\pi=0.124$). Further, project visibility was the third-most listed benefit and ranked second in significance ($L=13$, $\pi=0.142$) respective to other characteristics discovered, reflecting an appreciation towards exposure and openness to criticism. Alomar et al. [15] also find that the usage of bug bounty programs indicates higher security posture and how wide project visibility can help discover vulnerabilities that were previously not noticed by developers.
Interviewees expressed that having “many eyes” on the project and involving the community in searching for vulnerabilities ($I=9$) allows vulnerability detection that is otherwise impossible for OSS project maintainers alone. In particular, one interviewee stated, “When we implement something and maintain it for a number of years, we try to think of all possible edge cases and put ourselves in the shoes of an attacker trying to secure stuff; but it’s inevitable that we miss some things” (P9). Others also mentioned how project exposure allowed older security bugs, some as old as ten years, to be discovered and patched ($I=3$). Interviewees also recognize the importance of fixing vulnerabilities causing a “ripple effect” in the software supply chain ($I=3$), e.g., alerting affected projects to upgrade.
On the other hand, others had received and reviewed reports that were deemed low-quality because of their minor contributions ($I=6$). For instance, one interviewee described how after applying a fix to a low-quality report they received, the “real world quality of the software has essentially not really been improved” (P11), leaving them feeling like their time was wasted and an established negative association with receiving bug bounty report notifications. Results regarding OSS project maintainers’ perspectives on low-quality and spam reports are presented in Section 5.3.1.